Paul is 3 times as old as anna, in 12 years , he will be twice as old as anna . What is pauls age?
Answer by fractalier(6550)

Always set the variables in age problems NOW.
Thus let P be Paul's age NOW, and A be Anna's NOW.
Thus
P = 3A
In 12 years,
P + 12 = 2(A + 12)
now substitute in for P and get
3A + 12 = 2(A + 12)
3A + 12 = 2A + 24
A + 12 = 24
A = 12
Then, back-substitute and get
P = 36